OpenStudy (anonymous):

Please help me with this standard form question! 6.The speed of light is approximately 3 x 10^8 metres per second. How far will light travel in: a) 10 seconds b) 20 seconds c) 102 seconds. I tried multiplying by ten for the first one and then I ran out of ideas.

ganeshie8 (ganeshie8):

Yes distance = speed*time

ganeshie8 (ganeshie8):

you just need to multiply

ganeshie8 (ganeshie8):

for a, distance = 3 x 10^8 m/s * 10 s = 3 x 10^9 m

ganeshie8 (ganeshie8):

see if you can try b

OpenStudy (jadedry):

You can convert the seconds into standard form: x * 10^y 10 = 1 * 10^1 20 = 2 *10^1 102 = 1.02 *10^2 When multiplying, multiply 3 by "x" and 10^8 by y. 3*x = 3x 10^8 * 10^y = 10^8+y Can you solve them now?

OpenStudy (anonymous):

Actually, still not a hundred percent. When multiplying wouldnt it be 30 x 10^18 ? Can you please explain why you dont multiply the three and why the exponent only increased by one?

OpenStudy (unklerhaukus):

d = 3 x 10^8 m/s x 10 s = 3 x 10^8 x 10 m = 3 x 10^8 x 10^1 m = 3 x 10^{8+1} m = 3 x 10^9 m

OpenStudy (anonymous):

Thank you. :)